### v2

@@ Abstract
It is shown that condensation and precipitation do not disturb the hydrostatic equilibrium if the local pressure sink (condensation rate expressed in pressure units) is proportional to the local pressure, with a proportionality coefficient $k$ that is independent of altitude. In the real atmosphere, condensation rate is controlled, among other factors, by the vertical velocity that can vary freely over height. This means that, in general, condensation disturbs hydrostatic equilibrium and thus causes pressure adjustments through redistribution of air masses. It is proposed that $k$ maximised in the upper atmosphere results in additional upward motion, which leads to cyclone strengthening. Conversely, $k$ maximised closer to the surface produces additional downward motion, which causes cyclone’s weakening. The maximum scale of both effects should be set by the strength of the mass sink (precipitation). Using observational data, it is found that the mean intensification and weakening rates ($8$ and $6$ hPa day-1, respectively) in Atlantic tropical cyclones constitute about two thirds of their maximum concurrent precipitation (multiplied by gravity). This means that the storms intensify at positive vertically integrated air convergence – a pattern impossible in modeled dry hurricanes – with the pressure tendency being negative solely due to the fact that precipitation exceeds the vertically integrated moisture convergence by absolute magnitude. The implications of these results for recent studies evaluating the (de-)intensification process based on a mass continuity equation that neglects the mass sink are discussed.
@@ 1 Introduction
In a hydrostatic atmosphere, surface pressure reflects the weight of the atmospheric column. Consequently, the surface pressure tendency indicates changes in that column’s mass. According to the law of mass conservation, the rate of change of air mass in the column is equal to the vertically integrated air convergence plus evaporation $E$ minus precipitation $P$. This leads to the following equation for surface pressure tendency:
[…]
as derived, for example, in Trenberth (1991, Eq. (5)) and Wacker et al. (2006, Eq. (23)). Here, $p$ is surface pressure, $g$ is the acceleration of gravity, $\rho$ is the mass density of air, $\mathbf{u}$ is the horizontal velocity, and $H$ is a height at which $\rho$ can be assumed negligible,
@@
[…]
as in Eq. (6) of Lackmann and Yablonsky (2004). Here, $u_{r}$ denotes the radial velocity, and the averaging is performed over the cylinder’s area. Evaporation is neglected, as it is generally small in the storm core (see, e.g., Makarieva et al., 2017, their Table1).
Lackmann and Yablonsky (2004), using a numerical model, found that during the intensification phase of Hurricane Lili (2002), the average pressure drop within a $r=100$ km radius from the storm center was approximately $\partial\overline{p}/\partial t=-0.5$ hPa/hour. At the same time, the average pressure-equivalent mass loss from precipitation was three times greater: $-g\overline{P}=-1.5$ hPa/hour. This led Lackmann and Yablonsky (2004) to conclude that “the amount of atmospheric mass removed via precipitation exceeded that needed to explain the model sea level pressure decrease”.
@@
$-g\overline{P}=k\frac{\partial\overline{p}}{\partial t}$ (3)
with $k\sim 3$, omitting the precipitation term from the pressure tendency equation (2) and diagnosing pressure change solely from the air convergence term (the first term on the right-hand side of Eq. (2)) would result in a substantial error: the predicted pressure tendency would not only be twice as large in absolute magnitude as the observed value but also incorrect in sign (positive). In other words, neglecting precipitation would suggest a rapidly weakening storm instead of a rapidly intensifying one.
We emphasize that $k>1$ in Eq. (3) implies that the storm intensifies with a positive vertically integrated air convergence (i.e., more air flows laterally into the cylinder than out). This stands in fundamental contrast to dry circulations, where a surface pressure drop can occur only if the vertically integrated air convergence is negative.
Wacker et al. (2006) reported a similar pattern in a modeled cyclone that deepened by about $60$ hPa over 90 hours (a mean rate of $0.7$ hPa/hour), while the mean precipitation rate in the eyewall over six hours was around $3.7$ hPa/hour – a value the authors described as “a remarkable pressure drop”.
This again implies that the convergence term was positive, and confirms that neglecting the precipitation term in the pressure tendency equation can lead to major inaccuracies.
@@ 2 The model of Sparks and Toumi
3 Methods
To analyze the dependence between intensification rate and precipitation we followed the approach of Makarieva et al. (2017). We used the EBTRK dataset released on 27 July 2016 (Demuth et al., 2006) and the 3-hourly TRMM 3B42 (version 7). EBTRK data are recorded every six hours. For the years 19981998199819982015201520152015, for each $k$-th record in the EBTRK dataset (with $(k-1)$th and $(k+1)$th records referring to the same storm), we defined the intensification rate $I_{k}\equiv-2(p_{k+1}-p_{k-1})$ (hPa day-1), where $p$ (hPa) is the minimum central pressure. We then selected all tropical storms over land (those with a negative value of distance to land in the last column of the EBTRK file) with the known radius of maximum wind: a total of $360$ values of $I$, of which $19$ were zero, $301$ negative (weakening storms) and $40$ positive (intensifying storms).
Using the TRMM data (spatial resolution $0.25^{\rm o}$ latitude $\times$ $0.25^{\rm o}$ longitude), for each $k$-th position of the storm center in EBTRK, we established the dependence of precipitation $P(r)$ on distance $r$ from the storm center, with $P(r_{i})$ defined as the mean precipitation in all grid cells with $r_{i}<25i$ km, $1\leq i\leq 120$, $r_{i}\equiv 25(i-1)+12.5$ km. Examples of $\overline{P}(r)$ distributions for individual hurricanes are given in Fig. 11 of Makarieva et al. (2017). Maximum value of thus obtained $\overline{P}(r_{i})$ and radius $r_{i}$ corresponding to this maximum were defined as $P_{m}\equiv\overline{P}(r_{P})$ and $r_{P}$, respectively, for the $k$-th record. Additionally, $\overline{P}(r_{i})$ for which $25(i-1)\leq r_{m}<25i$ km was defined as precipitation $\overline{P}(r_{m})$ at the radius of maximum wind $r_{m}$ for storms where $r_{m}$ was known. To enable numerical comparison between precipitation and intensification rates [^1] For example, one mm of water per hour (multiplied by $\rho_{l}g$, where $\rho_{l}$ is the density of liquid water) is equivalent to $2.4$ hPa per day., we expressed precipitation in hPa day-1 by multiplying precipitation by factor $g$.
Figure 1: Intensification rates taken by absolute value $|I|$ and maximum precipitation $gP_{m}$ in intensifying and weakening storms on land and over the ocean (shown for comparison). Numbers of storms in each group are shown along the lower horizontal axis. Medians of $I$ and $gP_{m}$ are shown along the upper horizontal axis. Note the logarithmic scale on the vertical axis. Crosses show mean values.
@@ 4 Results
The median, lower and upper quartiles for storms weakening over land are $I=-14$ ($-30$, $-5.5$) hPa day-1, while their maximum concurrent precipitation is $gP_{m}=7.7$ ($4$, $14$) hPa day-1. For storms intensifying over land $I=4$ ($2$, $6$) hPa day-1 and $gP_{m}=9.1$ ($3$, $17$) hPa day-1 (Fig. 1).
Intensification rates shown in Fig. 1 describe changes of the minimum surface pressure and thus correspond to the central pressure tendency in the axisymmetric model of Sparks and Toumi (2022a). Minimum surface pressure by definition changes faster than the mean surface pressure. Thus applying Eq. (6) to the circle $r\leq r_{P}$ and neglecting the precipitation term in Eq. (6) should overestimate the absolute magnitude of $\partial\overline{p}/\partial t$ in the weakening storms by at least $gP_{m}/|I|\times 100\%\simeq 50\%$ and underestimate it in intensifying storms by $gP_{m}/|I|\times 100\%=220\%$. For storms weakening and intensifying over the ocean, the corresponding inaccuracies would be remarkably similar at $140\%$ and $150\%$ (Fig. 1).
Figure 2: Radius of maximum precipitation $r_{P}$ and radius of maximum wind $r_{m}$ in storms where $r_{m}$ is known. Numbers of storms in each group are the same as in Fig. 1. Medians of $r_{P}$ and $r_{m}$ are shown along the upper horizontal axis. Note the logarithmic scale on the vertical axis. Crosses show mean values.
Figure 2 shows that the radius of maximum precipitation $r_{P}$ is somewhat larger than the radius of maximum wind $r_{m}$ that Sparks and Toumi (2022a) considered in their model, Eq. (7). Precipitation within the radius of maximum wind $\overline{P}(r_{m})$ is, by definition, lower than $P_{m}$. It was $g\overline{P}(r_{m})=5.5$ ($2$, $11$) hPa day-1 for weakening and $g\overline{P}(r_{m})=6.2$ ($2$, $12$) hPa day-1 for intensifying storms over land. The median, lower and upper quartiles for the ratio $g\overline{P}(r_{m})/|I|\times 100\%$ were $40$ ($13$, $100$)for weakening and $160$ ($50$, $290$)for intensifying storms. This means that in both weakening and intensifying storms the vertically integrated air convergence is generally positive, such that the sign of the pressure tendency is decided by the absolute magnitude of precipitation rate.
@@
$-2p(r)\frac{\mathcal{U}_{s}(r)}{r}\equiv\frac{\partial\overline{p}}{\partial t},$ (8)
@@
The mean column speed, $\chi$, as acknowledged by Sparks and Toumi (2022a), is a non-observable quantity, with its absolute magnitude being at least two orders of magnitude smaller than the actual radial velocities associated with inflow and outflow. In contrast, precipitation can be directly retrieved from observations. Sparks and Toumi (2022a) suggested that their model could be extended to tropical cyclones over the ocean. For such cases, the similarity between $|I|$ and $gP_{m}$ is even more pronounced (see Fig. 1), which led us to propose that precipitation may play a central role in both intensification and de-intensification of storms Makarieva and Nefiodov (2024). Given that precipitation is expected to increase as the radius of maximum wind contracts and vertical velocities intensify, this mechanism could explain the dependence of intensification and weakening rates on storm radius reported by Sparks and Toumi (2022b) using their model. We argue that focusing on surface pressure tendency offers a promising framework for understanding storm dynamics, provided that the role of the precipitation-induced mass sink is fully and accurately incorporated.
@@ 5 Pressure adjustments related to precipitation
$-\frac{\partial p}{\partial z}=\rho g=\frac{p}{h},\quad h\equiv\frac{RT}{Mg},$ (10)
@@
This means that the removal of condensate from the atmospheric column will not disturb the hydrostatic equilibrium if the pressure sink is proportional to pressure itself, i.e., if $\dot{p}(z)=kp(z)$, where $k$ is independent of altitude $z$.
@@
If $k(z)$ has a maximum, and condensation is predominantly concentrated, in the upper atmosphere (Fig. 3e), then the pressure adjustment will be predominantly directed upward leading to intensification (could convective bursts preceding rapid intensification be an example?) If, on the contrary, $k(z)$ is maximized, and condensation more intense, in the lower atmosphere, then the pressure adjustment will proceed downward suppressing the upwelling and possibly resulting in de-intensification (Fig. 3f). The intensification rate can become zero, and the storm steady, at some intermediate position of the condensation maximum [^3] With vertical velocity freely varying along height, these conclusions do not depend on our neglect of the difference in the molar masses of water vapor and dry air made when deriving Eq. (13). Whatever the exact form of Eq. (13), it does not contain vertical velocity.. In the view of $|I|\simeq gP$, the concepts shown in Fig. 3 are clearly relevant to storm intensification, but largely remain unstudied.
If, as indicated by the data in Fig. 1, the vertically integrated air convergence is usually positive, and the negative sign of the pressure tendency in the intensifying storms is ensured by precipitation, it is clear that relatively minor changes in precipitation can turn a strengthening storm into a weakening one. This pattern was observed by Su et al. (2020) who found that in all intensity categories, from tropical depressions to intense hurricanes, intensifying storms differ from weakening ones by a nearly universal precipitation surplus of 2 mm/hour in the storm core. Smith et al. (2021) observed, in a modeling study, that the appearance of a low-level outflow is associated with storm weakening. This is also understandable in terms of the precipitation impact: unlike a high-level outflow, a low-level one does not generate much precipitation. So as soon as the air ceases to ascend, precipitation drops and the storm begins to weaken.
@@
We argue that focusing on surface pressure tendency offers a promising framework for understanding storm intensification and weakening, provided that the role of the precipitation-induced mass sink is fully and accurately incorporated.

### v3

@@ Abstract
It is shown that condensation and precipitation do not disturb hydrostatic equilibrium if the local pressure sink (the condensation rate expressed in pressure units) is proportional to the local pressure, with a proportionality coefficient $k$ that is independent of altitude. In the real atmosphere, however, the condensation rate depends, among other factors, on the vertical velocity, which varies with height. As a result, condensation generally disturbs hydrostatic equilibrium and induces pressure adjustments through air-mass redistribution. We propose that a profile in which $k$ is maximized in the upper atmosphere leads to additional upward motion and cyclone intensification, whereas a maximum closer to the surface induces downward motion and cyclone weakening. The magnitude of both effects is expected to be set by the strength of the precipitation mass sink. Using observational data, we find that the median intensification and weakening rates—$12$ and $8$ hPa day-1, respectively, measured over six-hour intervals in Atlantic tropical cyclones—amount to about three quarters of the maximum concurrent precipitation rate (multiplied by gravity) in the core precipitation region. This implies intensification under conditions of positive vertically integrated air convergence, a regime impossible in modeled dry hurricanes, with the negative pressure tendency arising because precipitation exceeds the vertically integrated moisture convergence by absolute magnitude. The implications of these results for recent studies that evaluate tropical cyclone (de-)intensification using mass continuity equations that neglect the precipitation mass sink are discussed.
@@ 1 Introduction
In a hydrostatic atmosphere, surface pressure reflects the weight of the atmospheric column. Consequently, the surface pressure tendency indicates the rate at which that column’s mass change. According to the law of mass conservation, the rate of change of air mass in the column is equal to the vertically integrated air convergence plus evaporation $E$ minus precipitation $P$. This leads to the following equation for surface pressure tendency:
$\frac{\partial p}{\partial t}=-g\int_{0}^{H}\mathrm{div}(\rho\mathbf{u})dz+g(E-P),$ (1)
as derived, for example, in Trenberth (1991, Eq. (5)) and Wacker et al. (2006, Eq. (23)). Here, $p$ is surface pressure, $g$ is the acceleration of gravity, $\rho$ is the mass density of air, $\mathbf{u}$ is the horizontal velocity, and $H$ is a height at which $\rho$ can be assumed negligible.
@@
$\frac{\partial\overline{p}}{\partial t}=-\frac{2g}{r}\int\limits_{0}^{H}\rho u_{r}dz-g\overline{P},$ (2)
as in Eq. (6) of Lackmann and Yablonsky (2004). Here, $u_{r}$ denotes the radial velocity, and the averaging is performed over the cylinder’s base area. Evaporation is neglected, as it is generally small in the storm core (see, e.g., Makarieva et al., 2017, their Table1).
Using a numerical model, Lackmann and Yablonsky (2004) found that during the intensification phase of Hurricane Lili (2002), the average pressure drop within a radius of $r=100$ km from the storm center was approximately $\partial\overline{p}/\partial t=-0.5$ hPa h-1. Over the same period, the average pressure-equivalent mass loss due to precipitation was about three times larger, with $g\overline{P}=-1.5$ hPa h-1. On this basis, Lackmann and Yablonsky (2004) concluded that “the amount of atmospheric mass removed via precipitation exceeded that needed to explain the model sea level pressure decrease.”
@@
$g\overline{P}=c\frac{\partial\overline{p}}{\partial t}$ (3)
with $c\sim 3$, omitting the precipitation term from the pressure tendency Eq. (2) and diagnosing pressure change solely from the air convergence term (the first term on the right-hand side of Eq. (2)) would result in a substantial error: the predicted pressure tendency would not only be twice as large in absolute magnitude as the observed value but also incorrect in sign (positive). In other words, neglecting precipitation would suggest a rapidly weakening storm instead of a rapidly intensifying one.
We emphasize that $c>1$ in Eq. (3) implies that the storm intensifies with a positive vertically integrated air convergence (i.e., more air flows laterally into the cylinder than out). This stands in fundamental contrast to dry circulations, where a surface pressure drop can occur only if the vertically integrated air convergence is negative. From a dynamical perspective, a value of $c>1$ implies that, in the upper troposphere, the combined pressure-gradient and centrifugal forces are insufficient to export from the storm core as much air as is drawn inward by the low-level pressure gradient.
Wacker et al. (2006) reported a similar pattern in a modeled cyclone that deepened by about $60$ hPa over 90 hours (a mean rate of $0.7$ hPa h-1), while the mean precipitation rate in the eyewall over six hours was approximately $3.7$ hPa h-1—a value the authors described as “a remarkable pressure drop”. This again implies a positive convergence term and confirms that neglecting the precipitation term in the pressure tendency equation can lead to substantial inaccuracies.
@@ 2 The model of Sparks and Toumi
To analyze the relationship between intensification rate and precipitation, we followed Makarieva et al. (2017). We used the EBTRK dataset (released 27 July 2016; Demuth et al., 2006) together with the 3-hourly TRMM 3B42 precipitation product (version 7). EBTRK records are available at 6-hour intervals.
For 1998–2015, and for each $k$th EBTRK record (with the $(k+1)$th record referring to the same storm), we defined the intensification rate as $I_{k}\equiv-4(p_{k+1}-p_{k})$ (hPa day-1), where $p$ is the minimum central pressure (hPa). We selected tropical storms over land, identified by a negative distance-to-land value in the final EBTRK column, for which the radius of maximum wind is available. This yielded 360 values of the intensification rate $I$: 45 equal to zero, 277 negative (weakening), and 38 positive (intensifying).
Using the TRMM data (spatial resolution $0.25^{\circ}$ latitude $\times$ $0.25^{\circ}$ longitude), for each $k$th position of the storm center in EBTRK, we established the dependence of precipitation on distance $r$ from the storm center, with $\overline{P}_{k}(r_{i})$ defined as the mean precipitation in all grid cells with $r_{i}<25i$ km, $1\leq i\leq 120$, $r_{i}\equiv 25(i-1)+12.5$ km. Examples of precipitation distributions for individual hurricanes are given in Fig. 11 of Makarieva et al. (2017).
The maximum value of thus obtained $\overline{P}_{k}(r_{i})$ and the radius $r_{i}$ corresponding to this maximum were defined as $\overline{P}_{k}(r_{P})$ and $r_{P}$, respectively, for the $k$th record. Additionally, $\overline{P}_{k}(r_{i})$ for which $25(i-1)\leq r_{m}<25i$ km was defined as precipitation $\overline{P}_{k}(r_{m})$ at the radius of maximum wind $r_{m}$ for storms where $r_{m}$ was known. To enable numerical comparison between precipitation and intensification rates [^1] For example, one mm of water per hour (multiplied by $\rho_{l}g$, where $\rho_{l}$ is the density of liquid water) is equivalent to $2.4$ hPa per day., we expressed precipitation in hPa day-1.
To partially account for the different temporal resolutions of EBTRK (6 hours) and TRMM (3 hours), we compared $I_{k}$ with $P_{m}\equiv[\overline{P}_{k}(r_{P})+\overline{P}_{k+1}(r_{P})]/2$ and $P_{r}\equiv[\overline{P}_{k}(r_{m})+\overline{P}_{k+1}(r_{m})]/2$. However, the results were found to be largely insensitive to extending the time scale used to define the intensification rate to 12 hours, defining it as $I_{k}\equiv-2(p_{k+1}-p_{k-1})$, while comparing it with the 3-hour precipitation at each $k$th position of the storm. In all cases, precipitation was found to be similar in magnitude to the intensification rate.
Figure 1: Intensification rates taken by absolute value $|I|$ and maximum precipitation $gP_{m}$ in intensifying and weakening storms on land and over the ocean (shown for comparison). Numbers of storms in each group are shown along the lower horizontal axis. Medians of II and $gP_{m}$ are shown along the upper horizontal axis. Note the logarithmic scale on the vertical axis. Crosses and dots show mean values and outliers, respectively.
@@ 3 Results
The median, lower and upper quartiles for storms weakening over land are I=16I=-16 (32-32, 8-8) hPa day-1, while their maximum concurrent precipitation is $gP_{m}=8$ (44, $14$) hPa day-1. For storms intensifying over land $I=4$ (44, 88) hPa day-1 and $gP_{m}=11$ (55, $14$) hPa day-1 (Fig. 1).
Intensification rates shown in Fig. 1 describe changes of the minimum surface pressure and thus correspond to the central pressure tendency in the axisymmetric model of Sparks and Toumi (2022a). Minimum surface pressure by definition changes faster than the mean surface pressure. Thus applying Eq. (6) to the circle $r\leq r_{P}$ and neglecting the precipitation term in Eq. (6) should overestimate the absolute magnitude of $\partial\overline{p}/\partial t$ by at least $gP_{m}/|I|\times 100\%\simeq 50\%$ in the weakening storms and by $275\%$ in the intensifying storms on land. For storms weakening and intensifying over the ocean, the corresponding inaccuracies would be remarkably similar at $130\%$ and $120\%$ (Fig. 1). According to Eq. (2), this implies that, in both weakening and intensifying storms, the vertically integrated air convergence is generally positive, such that the sign of the pressure tendency is determined by the absolute magnitude of the precipitation rate.
Figure 2: Radius of maximum precipitation $r_{P}$ and radius of maximum wind $r_{m}$ in storms where $r_{m}$ is known. Numbers of storms in each group are the same as in Fig. 1. Medians of $r_{P}$ and $r_{m}$ are shown along the upper horizontal axis. Note the logarithmic scale on the vertical axis. Crosses and dots show mean values and outliers, respectively.
Figure 2 shows that the radius of maximum precipitation, $r_{P}$, is somewhat larger than the radius of maximum wind, $r_{m}$, considered by Sparks and Toumi (2022a) in their model [Eq. (7)]. By definition, precipitation within the radius of maximum wind, $\overline{P}_{k}(r_{m})$, is lower than $\overline{P}_{k}(r_{P})$. For storms over land, the median (lower, upper) quartiles of $gP_{r}$ were $4.9$ ($2.6$, $10$) hPa day-1 for weakening storms and 66 ($3.3$, $10$) hPa day-1 for intensifying storms. The median (lower, upper) quartiles of $gP_{r}/|I|\times 100\%$ were $32$ ($13$, 6969)% and $120$ ($36$, 220220)% for weakening and intensifying storms, respectively.
@@
$\frac{\partial\overline{p}}{\partial t}\equiv-2p(r)\frac{\mathcal{U}_{s}(r)}{r},$ (8)
@@
The mean column speed, $\chi$, as acknowledged by Sparks and Toumi (2022a), is a non-observable quantity whose absolute magnitude is at least two orders of magnitude smaller than the actual radial velocities associated with inflow and outflow. In contrast, precipitation can be directly retrieved from observations. Sparks and Toumi (2022a) further suggested that their model could be extended to tropical cyclones over the ocean. For such cases, the similarity between $|I|$ and $gP_{m}$ is even more pronounced (see Fig. 1), which led us to propose that precipitation may play a central role in both the intensification and de-intensification of storms (Makarieva and Nefiodov, 2024). Because precipitation is expected to increase as the radius of maximum wind contracts and vertical velocities intensify, this mechanism could also explain the dependence of intensification and weakening rates on storm radius reported by Sparks and Toumi (2022b) using their model.
@@ 4 Pressure adjustments related to precipitation
$\frac{\partial p}{\partial z}=\rho g=\frac{p}{h},\qquad h\equiv\frac{RT}{Mg},$ (10)
@@
This means that the removal of condensate from the atmospheric column will not disturb the hydrostatic equilibrium if the pressure sink is proportional to pressure itself, i.e., if
$\dot{p}(z)=kp(z),$ (14)
where $k$ is independent of altitude $z$.
@@
If $k(z)$ has a maximum and condensation is predominantly concentrated in the upper atmosphere (Fig. 3e), the resulting pressure adjustment will be directed mainly upward, leading to storm intensification (convective bursts preceding rapid intensification may be an example of this regime). If, by contrast, $k(z)$ is maximized and condensation is more intense in the lower atmosphere, the pressure adjustment will proceed downward, suppressing upwelling and potentially resulting in storm de-intensification (Fig. 3f). At some intermediate height of the condensation maximum, the intensification rate can become zero, corresponding to a steady storm state [^3] With vertical velocity freely varying with height, these conclusions do not depend on our neglect of the difference in the molar masses of water vapor and dry air made when deriving Eq. (13). Whatever the exact form of Eq. (13), it does not contain vertical velocity..
In view of $|I|\simeq gP$, the concepts illustrated in Fig. 3 are therefore directly relevant to storm intensification, yet they remain largely unexplored.
Figure 3 illustrates that, although the pressure tendency is determined by the precipitation mass sink, it pertains primarily to dry air, since the void created by water vapor condensation is filled by atmospheric air in which dry air dominates. This can also be seen by considering the mass budgets of atmospheric water and dry air separately, as suggested by a reviewer. Using the mass budget equation for water in Eq. (1),
$\frac{\partial M_{v}}{\partial t}+\int_{0}^{H}\mathrm{div}(\rho_{v}\mathbf{u})dz=E-P,$ (15)
where $M_{v}$ is the total mass of precipitable water in the atmospheric column, we obtain
$\frac{\partial p}{\partial t}-g\frac{\partial M_{v}}{\partial t}=-g\int_{0}^{H}\mathrm{div}(\rho_{d}\mathbf{u})dz.$ (16)
Since the surface pressure of an intensifying tropical cyclone typically decreases by several tens of millibars (Chavas et al., 2017), whereas the total precipitable water content increases by only about 20 mm (Makarieva et al., 2017, their Fig. 4a)—corresponding to a surface pressure increase of roughly 2 mb—the second term on the left-hand side of Eq. (16) can, in most circumstances, be neglected. This implies that the surface pressure tendency in tropical storms is, to high precision, determined by changes in the dry-air mass [^4] The mean column velocity may then be defined by Eq. (5), with $\rho=\rho_{d}+\rho_{v}$ replaced by the dry-air density $\rho_{d}$. Under this reformulation, the model of Sparks and Toumi (2022a) would be formally valid. However, although $\rho$ and $\rho_{d}$ differ by only about 1%, the column-mean velocities obtained using $\rho_{d}$ and those obtained using $\rho$ with the precipitation term neglected can differ by more than 100% and may even have opposite signs. It is therefore essential to specify the definition of air density in Eq. (5) precisely in order to ensure the internal consistency of the model. Because the mean column velocity is defined using $\rho$ in both Sparks and Toumi (2022a) and the companion paper by Sparks and Toumi (2022b), this requirement is not satisfied..
Taken together, the data in Fig. 1 and Eq. (16) indicate that, in intensifying storms, the tendency of dry-air mass in the atmospheric column is of the same order as the precipitation rate within the radius of maximum precipitation. In other words, within the core precipitation region of the storm, the dry-air content decreases at a rate comparable to that at which precipitation removes water from the column [^5] The central pressure tendency overestimates the mean pressure tendency; consequently, the relative contribution of precipitation to the mean pressure tendency is greater than indicated by the data in Fig. 1. Additionally, the coarse spatial resolution of TRMM 3B42 is expected to underestimate precipitation in the most intense storms with small radii of maximum wind..
We argue that this correspondence warrants further investigation. By removing water vapor from the column, precipitation perturbs the vertical pressure distribution and induces compensating air motions that would be absent in a dry atmosphere. Evidence for the dynamical importance of this precipitation-induced mass sink is provided by both observations and numerical models. Pressure gradients that drive atmospheric motion act equally on dry air and water vapor. The presence of positive vertically integrated air convergence in cyclones therefore implies that upper-level pressure gradients, together with centrifugal forces, are unable to export air from the storm core at a rate sufficient to offset the low-level moisture inflow. As a result, relatively small changes in precipitation rate can shift a cyclone from intensification to weakening. Consistent with this interpretation, Su et al. (2020) found that, across all intensity categories from tropical depressions to major hurricanes, intensifying storms exhibit a nearly universal precipitation surplus of approximately 2 mm h-1 in the storm core relative to weakening storms. Further support comes from modeling studies showing that suppressing precipitation fallout, while retaining latent heat release, strongly weakens storm dynamics (Bryan and Rotunno, 2009; Wang and Lin, 2021). In a modeling study, Smith et al. (2021) found that the development of a low-level outflow is associated with storm weakening. This behavior can also be interpreted in terms of the role of precipitation: unlike a high-level outflow, a low-level outflow is not conducive to substantial precipitation. As upward motion weakens, precipitation decreases, and the storm correspondingly weakens.
@@
At present, these patterns lack a quantitative theoretical explanation. We suggest that focusing on surface pressure tendency provides a promising framework for understanding tropical cyclone dynamics, provided that the precipitation-induced mass sink is incorporated fully and consistently.